# LLMs in the Heart of Differential Testing: A Case Study on a Medical Rule Engine

### v1

@@ Abstract.
The Cancer Registry of Norway (CRN) uses an automated cancer registration support system (CaReSS) to support core cancer registry activities, i.e, data capture, data curation, and producing data products and statistics for various stakeholders. GURI is a core component of CaReSS, which is responsible for validating incoming data with medical rules. Such medical rules are manually implemented by medical experts based on medical standards, regulations, and research. Since large language models (LLMs) have been trained on a large amount of public information, including these documents, they can be employed to generate tests for GURI. Thus, we propose an LLM-based test generation and differential testing approach (LLMeDiff) to test GURI. We experimented with four different LLMs, two medical rule engine implementations, and real medical rules to investigate the hallucination, success, time efficiency, and robustness of the LLMs to generate tests, and these tests’ ability to find potential issues in GURI. Our results showed that GPT-3.5 hallucinates the least, is the most successful, and is generally the most robust; however, it has the worst time efficiency. Our differential testing revealed 22 medical rules where implementation inconsistencies were discovered (e.g., regarding handling rule versions). Finally, we provide insights for practitioners and researchers based on the results.
1. Introduction
@@
The cancer registration support system (CaReSS) is a real-world socio-technical software system developed and operated by the Cancer Registry of Norway (CRN), a public organization in Norway under the National Institute of Public Health (NIPH). CaReSS receives cancer patient data through cancer messages from other public and private entities, including hospitals, laboratories, and health registries (e.g., national patient registry). Such data is then processed by CaReSS to produce outputs (including data and statistics) for its end users, such as the public, researchers, and policymakers.
@@
One core functionality of CaReSS is to validate incoming cancer messages and medical data with a dedicated rule validation software system called GURI, which checks the correctness of received cancer messages, validates the data with various medical rules, aggregates individual cancer messages into cancer cases, and validates the aggregation. The rules are manually defined by medical coders based on various standards and regulations, such as International Classification of Diseases (ICD)-10 (10th edition), ICD-O-2 (ICD for Oncology, 2nd edition) for solid tumors, and ICD-O-3 (ICD for Oncology, 3rd edition) for non-solid tumors. Extensively testing GURI is important since an incorrect implementation would produce erroneous data or statistcs, which are used in medical research, monitoring and evaluation of cancer care.
@@
As large language models (LLMs) have been trained on a large amount of public information, including the standards, regulations, and other documents that are the sources for specifying the medical rules, we employ these LLMs to generate tests for these medical rules. Consequently, we developed an LLM-based differential testing approach called LLMeDiff to find potential faults in GURI. An LLM is used to generate tests from medical rules (that yield Pass, Fail, and NotApplied results for each rule), which are then executed on GURI and a simple reference implementation (Dvare++). A mismatch between the outputs from these two flags a potential inconsistency stemming from an implementation issue in GURI.
@@
In our previous works, we studied a state-of-the-art test generation tool (i.e., EvoMaster [3]) applied to GURI [26], reduced its number of executed tests with machine learning (ML) classifiers [21], and built cyber-cyber digital twins of GURI [29]. In contrast, in this paper, we improve on our previous works by (1) leveraging LLMs to generate valid medical data as test inputs, (2) addressing the oracle problem of rule results through differential testing with a reference implementation, and (3) reducing the number of executed tests against GURI by only generating three tests per rule.
@@
We experimented LLMeDiff by selecting four LLMs (i.e., Mistral, Llama2, Mixtral, and GPT-3.5), real medical rules and their mutations, and two rule engine implementations (i.e., GURI and Dvare++). We focused on two perspectives: (1) how effective, efficient, and robust the selected LLMs are in generating tests from medical rules; and (2) how many potential issues we can identify in the implementation of GURI with our differential testing approach.
@@
Our results showed that LLMs are highly effective in generating medical rule tests. In particular, GPT-3.5 is the most effective LLM with hardly any hallucinations (i.e., mean completion rate of %) and being successful in generating Pass (%), Fail (%), and NotApplied (%) tests. However, GPT-3.5 is also the least efficient model with compared to the fastest models, i.e., Mixtral with and Mistral with . In terms of robustness, GPT-3.5 performs the best for Pass tests with % and equally to Mixtral and Mistral for NotApplied tests with %. Only for Fail tests, GPT-3.5 with % is inferior to Mistral and Mixtral with both having %. Finally, regarding the differential testing results, we observe that Mixtral has the highest number of matches, while Mistral has the highest number of mismatches per rule. A total of 44 rule mismatches between GURI results and Dvare++ for Mistral, followed by GPT-3.5 with 32, Llama2 with 29 and Mixtral with 27. While not all mismatches can be considered faults, our analysis reveals 22 rules that show inconsistencies in terms of handling either rule versions, date format, or variable dependencies.
These results suggest that LLMs are effective rule test generators to discover mismatches with differential testing, improving the current state of practice at the CRN. Whether a mismatch is an actual fault still requires manual root cause analysis by the CaReSS developers.
2. Real-World Application Context
@@
Dvare111https://github.com/dvare/dvare-framework/ is a rule-expression language engine that runs the rules on the data and asserts the results into true or false depending on if any of the condition matches. GURI employs Dvare as the underlying engine to evaluate cancer messages against medical rules. While Dvare yields a Boolean result, GURI categorizes the results based on logical operations, particularly the implies operator. Consider an example rule with two operands joined by an implies operation, as illustrated in Fig. 1. In this scenario, GURI handles the rules as follows: (1) If the LeftOperand is true, the rule is applied and can be either Pass or Fail depending on the RightOperand (true for Pass, and false for Fail). (2) If the LeftOperand is false, the rule is NotApplied, irrespective of the RightOperand. This categorization enables GURI to present rule validation outcomes with enhanced clarity for analysis and prediction within CaReSS.
Testing GURI is essential to ensure the accuracy of the cancer data and statistics produced by CaReSS [21]. The challenges of testing GURI are as follows. (1) Evolution: GURI and Dvare continuously evolve in response to changing cancer messages format and medical rules dictated by standards/regulations or through cancer research. (2) High-cost: Testing multiple evolving implementations of GURI with various underlying Dvare versions incurs high time and resource costs. (3) Labor-intensive: Manual testing of each version GURI in different development phases is a laborious task. Therefore, a cost-effective and automated approach is necessary to test the evolving GURI [27]. (4) Absence of Test Oracles: For medical rule testing, there are no precise test oracles to determine whether the rules are implemented correctly, posing additional challenges for automated testing.
3. Approach
@@
fnum@@desciitemStep 1: be precise;
fnum@@desciitemStep 2: identify rule variables and generate tests:
@@
fnum@@desciitemStep 3: do not be repetitive;
fnum@@desciitemStep 4: express your confidence (confidence_score);
@@
fnum@@desciitemStep 5: generate a JavaScript Object Notation (JSON) object as the output.
@@ 4.2. LLM Selection and Settings
GPT-3.5 from OpenAI, is a generative pre-trained transformer, while Llama2, from Meta, is a pre-trained and fine-tuned LLM with 13 billion parameters. We choose Mistral from Mistral AI for its good performance in reasoning, mathematics, and code generation, outperforming Llama2 on multiple benchmarks [23]. Furthermore, Mixtral, a sparse mixture of experts model with open weights, distinguishes itself by matching or surpassing GPT-3.5 on various tasks [24]. Its impressive performance and widespread popularity within the AI community, particularly on platforms like HuggingFace222https://huggingface.co/ were the reasons to select it.
@@
(2) the OpenAI API reference333https://platform.openai.com/docs/api-reference/introduction/ defaults to 0.7;
(3) open-source frameworks such as The Rasa Community444https://rasa.com/docs/rasa/next/llms/llm-intent/ use 0.7;
@@
Regarding the prompt (see Section 3), only Llama2 and GPT-3.5 support separate system and user parts.
@@
The three models Mistral, Llama2, and Mixtral are accessed through the AI platform fireworks.ai555https://fireworks.ai/, while GPT-3.5 which is accessible only on the OpenAI platform666https://openai.com/.
4.3. Medical Rules and Rule Mutations
@@
In addition, to enlarge the studied rule set in RQ1 and compute the robustness metrics in RQ2, we define a set of mutation operators on medical rules (see Table 1), inspired by previous research [36, 22].
@@
In total, we retrieve mutated rules from the rules.
Note that we use the mutated rules only to evaluate the tests generated by the LLMs in RQ1 and RQ2 and not in RQ3, as GURI does not have these rules implemented and, hence, would not return the expected test results.
Table 1. Mutation operators
@@
4.4. Evaluation Metrics
In this section, we define the metrics used to answer the RQs.
4.4.1. RQ1
@@
4.4.2. RQ2
We assess the robustness of an LLM by introducing minor changes to each rule and study how much the outputs of the original and mutated rules differ.
@@ 4.5. Statistical Analyses
negligible: ,
small: ,
medium: ,
@@ 4.6. Threats to Validity
Internal validity is about the experiment setup.
We use hosted LLMs and do not run the models ourselves, which removes our control over how the models are run.
For example, we can not be sure that model instances and sessions are not shared across users and HTTP requests, and, therefore, information from previous prompts may leak into our results.
However, we studied the platforms’ documentation and are quite certain that each request is treated as its own session and that no leakage should happen.
Another threat is the LLM temperature setting, where we picked a common value and kept it consistent across the LLMs.
The results would likely change with different temperatures.
Moreover, as LLMs are stochastic in nature, our results might also suffer from random artifacts.
To address this threat, we follow the best practice from search-based software engineering and repeat our experiment times and use appropriate statistical tests [4].
Using different statistical procedures to draw conclusions might lead to slightly different results.
Finally, every piece of software is prone to bugs; so is our experimentation pipeline.
We carefully reviewed our scripts and results and reran the experiments multiple times to ensure we do not experience critical bugs.
@@
Finally, we perform a laboratory experiment, i.e., we run the tests against a standalone version of GURI in an isolated setting, and do not perform a field experiment, where the tests are executed against the real GURI.
5. Results and Analyses
In this section, we discuss the results for the RQs.
5.1. RQ1: Effectiveness and Efficiency
@@
Figure 3. Completion rates for all the rules per LLM.
Table 2. Statistical tests of the completion rates, success indices per test type, and inference times per LLM
@@
The statistical tests, depicted in Table 2, support these observations: GPT-3.5 is statistically better than Llama2 with medium and Mixtral and Mistral with large effect sizes.
Figure 4. Success indices for all the rules per test type and LLM.
@@
This shows that the LLMs struggle more to generate Fail tests, which might be due to the structure of the rules, where the LLM must satisfy the left clause of the implication while it must falsify the right clause (see Section 2).
Figure 5. Inference times for all the rules per LLM.
@@
5.2. RQ2: Robustness
Figure 6. Robustness indices for all the rules per test type and LLM.
Table 3. Statistical tests of the robustness indices for all the rules per test type and LLM
@@ 5.2. RQ2: Robustness
This is confirmed by the statistical tests (see Table 3): GPT-3.5 is better than Mixtral and Llama2 with medium and Mistral with large effect size.
@@
Finally for , we do not observe statistical differences between Mistral, Mixtral, and GPT-3.5 with %, %, and %, respectively.
@@ 5.3. RQ3: Differential Testing
A Warning result occurs when the rule execution is successful, but the data exhibits potential issues. For example, a patient’s age might be 120 years, which is theoretically possible but highly unlikely; hence the Warning result.
In addition, we identify two other result types:
@@
Table 4 shows all the observed matches and mismatches after executing each generated test.
@@
(1) Inconsistency in handling rule versions.We notice that some of the rules have both versions “active” (e.g., V09/1 and V09/2), while some other rules are “inactive” for the first version. Thus, such rules never Pass.
@@
(3) Inconsistency in handling variable dependencies.We notice that different variables are handled differently in terms of execution. For instance, the “ds” variable affects the whole cancer message (resulting in an empty response), while other variables (e.g., diagnose date) affect only specific rules (resulting always in NotApplied).
Table 4. Matches and mismatches between GURI and Dvare++ results, considering all 58 rules
@@
6. Discussion and Lessons Learned
In this section, we discuss our learned lessons.
6.1. Hallucination Effect
As an indicator of potential hallucination, we consider the LLMs accuracy in following the given instructions in the prompt.
We utilize the metric as a binary evaluation to determine whether the tests generated by the LLMs are correct.

### v4

@@ Abstract
The Cancer Registry of Norway (CRN) uses an automated cancer registration support system (CaReSS) to support core cancer registry activities, i.e., data capture, data curation, and producing data products and statistics for various stakeholders. GURI is a core component of CaReSS, which is responsible for validating incoming data with medical rules. Such medical rules are manually implemented by medical experts based on medical standards, regulations, and research. Since large language models (LLMs) have been trained on a large amount of public information, including these documents, they can be employed to generate tests for GURI. Thus, we propose an LLM-based test generation and differential testing approach (LLMeDiff) to test GURI. We experimented with four different LLMs, two medical rule engine implementations, and real medical rules to investigate the hallucination, success, time efficiency, and robustness of the LLMs to generate tests, and these tests’ ability to find potential issues in GURI. Our results showed that GPT-3.5 hallucinates the least, is the most successful, and is generally the most robust; however, it has the worst time efficiency. Our differential testing revealed 22 medical rules where implementation inconsistencies were discovered (e.g., regarding handling rule versions). Finally, we provide insights for practitioners and researchers based on the results.
Index Terms:
Large Language Models, Medical Rules, Automated Software Testing, Test Generation, Differential Testing, Electronic Health RecordsI Introduction
@@
The cancer registration support system (CaReSS) is a real-world sociotechnical software system developed and operated by the Cancer Registry of Norway (CRN), a public organization in Norway under the National Institute of Public Health (NIPH). CaReSS receives cancer patient data through cancer messages from, e.g., hospitals, laboratories, and other health registries. It then processes the incoming data to produce specialized data and statistics for its stakeholders, such as the public, researchers, and policymakers.
@@
One core functionality of CaReSS is to validate incoming cancer messages and medical data with a dedicated rule validation software system called GURI, which checks the correctness of received cancer messages, validates the data with medical rules, aggregates individual cancer messages into cancer cases, and validates the aggregation. The rules are manually defined by medical coders based on standards and regulations, such as International Classification of Diseases (ICD)-10 (10th edition), ICD-O-2 (ICD for Oncology, 2nd edition) for solid tumors, and ICD-O-3 (ICD for Oncology, 3rd edition) for non-solid tumors. Extensively testing GURI is important since an incorrect implementation would result in erroneous data or statistics, which are used in medical research as well as monitoring and evaluation of cancer care.
@@
As large language models (LLMs) are trained on a vast amount of public information, including the standards and regulations used for specifying CRN’s medical rules, we employ LLMs to generate tests from these rules in a differential testing approach called LLMeDiff to find potential faults in GURI. We assume that utilizing LLMs is advantageous over other methods (e.g., search or constraint solvers) due to their semantic knowledge of the medical context, ability to generate realistic and diverse inputs, and capability of understanding rule variable names in different languages (i.e., Norwegian in our case). LLMeDiff uses an LLM to generate tests from medical rules (that yield Pass, Fail, and NotApplied results for each rule), which are then executed on GURI and a simple reference implementation (Dvare++). A mismatch between the outputs potentially stems from an implementation issue in GURI.
@@
Previously, we studied a state-of-the-art test generation tool (i.e., EvoMaster [2]) applied to GURI [27], reduced its number of executed tests with machine learning (ML) classifiers [24], and built cyber-cyber digital twins of GURI [30]. In this paper, we (1) leverage LLMs to generate valid medical data as test inputs, (2) address the oracle problem of rule results through differential testing with a reference implementation, and (3) reduce the number of executed tests against GURI by only generating three tests per rule.
@@
We experimented with LLMeDiff by selecting four LLMs (i.e., Mistral, Llama2, Mixtral, and GPT-3.5), real medical rules and their mutations, and two rule engine implementations (i.e., GURI and Dvare++). We focused on two perspectives: (1) how effective, efficient, and robust the selected LLMs are in generating tests from medical rules; and (2) how many potential issues we can identify in the implementation of GURI with our differential testing approach.
@@
Our results showed that LLMs are highly effective in generating medical rule tests. In particular, GPT-3.5 is the most effective LLM with hardly any hallucinations (i.e., mean completion rate of %) and being successful in generating Pass (%), Fail (%), and NotApplied (%) tests. However, GPT-3.5 is also the least efficient model with compared to the fastest models, i.e., Mixtral with and Mistral with . In terms of robustness, GPT-3.5 performs the best for Pass tests with % and equally to Mixtral and Mistral for NotApplied tests with %. Only for Fail tests, GPT-3.5 with % is inferior to Mistral and Mixtral with both having %. Finally, regarding the differential testing results, we observe that Mixtral has the highest number of matches, while Mistral has the highest number of mismatches per rule. A total of 44 rule mismatches between GURI results and Dvare++ for Mistral, followed by GPT-3.5 with 32, Llama2 with 29 and Mixtral with 27. While not all mismatches can be considered faults, our analysis reveals 22 rules that show inconsistencies in terms of handling either rule versions, date format, or variable dependencies. These results suggest that LLMs are effective rule test generators to discover mismatches with differential testing, improving the current state of practice at the CRN. Whether a mismatch is an actual fault still requires manual root cause analysis by the CaReSS developers.
II Real-World Application Context
@@
Dvare111https://github.com/dvare/dvare-framework/ is a rule-expression language engine that runs the rules on the data and asserts the results into true or false depending on if any of the condition matches. GURI employs Dvare as the underlying engine to evaluate cancer messages against medical rules. While Dvare yields a Boolean result, GURI categorizes the results based on logical operations, particularly the implies operator. Consider an example rule with two operands joined by an implies operation, as illustrated in Fig. 1. This rule can be described as follows: If the surgical procedure (Kirurgi) is ’97’ and the basis code (Basis) is one of ’70’, ’75’, or ’76’, then the topography code (Topografi) must be ’619’ and the message type (Meldingstype) must be ’H’. In this scenario, GURI handles the rules as follows: (1) If the LeftOperand is true, the rule is applied and can be either Pass or Fail depending on the RightOperand (true for Pass, and false for Fail). (2) If the LeftOperand is false, the rule is NotApplied, irrespective of the RightOperand. This categorization enables GURI to present rule validation outcomes with enhanced clarity for analysis and prediction within CaReSS.
Testing GURI is essential to ensure the validity of the cancer data and statistics produced by CaReSS [24]. We identify four main challenges of testing GURI described in the following. (1) Evolution: GURI and Dvare continuously evolve in response to changing cancer message formats and medical rules (i.e., rule deletion, modification, or insertion) dictated by standards/regulations or through cancer research. (2) High-cost: Testing multiple evolving implementations of GURI with various underlying environments incurs high time and resource costs. (3) Labor-intensive: Manual testing of each version GURI in different development phases is a laborious task. Therefore, a cost-effective and automated approach is necessary to test the evolving GURI [28]. (4) Absence of Test Oracles: For medical rule testing, determining whether the rules are implemented correctly is largely manual, which presents significant challenges for automated testing. To address the absence of precise oracles, our approach uses differential testing to flag mismatches, supporting automated root cause analysis.
III Approach
@@
fnum@@desciitemStep 1 be precise;
fnum@@desciitemStep 2 identify rule variables and generate tests:
@@
fnum@@desciitemStep 3 do not be repetitive;
fnum@@desciitemStep 4 express your confidence (confidence_score);
@@
fnum@@desciitemStep 5 generate a JavaScript Object Notation (JSON) object as the output.
@@ IV-B LLM Selection and Settings
GPT-3.5 from OpenAI, is a generative pre-trained transformer, while Llama2, from Meta, is a pre-trained and fine-tuned LLM with 13 billion parameters.
We choose Mistral from Mistral AI for its good performance in reasoning, mathematics, and code generation, outperforming Llama2 on multiple benchmarks [13].
Furthermore, Mixtral, a sparse mixture of expert models with open weights, distinguishes itself by matching or surpassing GPT-3.5 on various tasks [14].
Its impressive performance and widespread popularity within the AI community, particularly on platforms like HuggingFace were the reasons to select it.
@@
(2) the OpenAI API reference defaults to 0.7;
(3) open-source frameworks such as The Rasa Community use 0.7;
@@
Regarding the prompt (see Section III), only Llama2 and GPT-3.5 support separate system and user parts.
@@
The three models Mistral, Llama2, and Mixtral are accessed through the AI platform fireworks.ai, while GPT-3.5 is accessible only on the OpenAI platform.
IV-C Medical Rules and Rule Mutations
@@
These rules exhibit varying levels of complexity across three aspects: clauses, variable names, and operators. The number of clauses ranges from 0 (representing rules with a single condition) to a maximum of 6. Similarly, the complexity of variable names varies from 1 (for simplistic rules with a single condition) to 7 (for the most complex rules), utilizing a total of 21 unique variable names. Furthermore, the employed operators range from 1 to 19, reflecting the diverse structural complexity of our rule dataset. For example, the rule in Fig. 1 has 4 variables (i.e., Kirurgi, Basis, Topografi, and Meldingstype), which are literals in a logical expression connected by logical operators (e.g., “and”, “or”, “not”). In this case, the “implies” operator connects the antecedent and the consequent clause. So, the correct count of clauses in this case is 2. As for the number of operators, we have 7 in total, including both logical and relational operators.
In addition, to enlarge the studied rule set in RQ1 and compute the robustness metrics in RQ2, we define a set of mutation operators on medical rules (see Table I), inspired by previous research [36, 25].
@@
In total, we retrieve mutated rules from the rules. Note that we use the mutated rules only to evaluate the tests generated by the LLMs in RQ1 and RQ2 and not in RQ3, as GURI does not have these rules implemented and, hence, would not return the expected test results.
Table I: Mutation operators
@@
IV-D Evaluation Metrics
IV-D1 RQ1
@@
IV-D2 RQ2
We assess the robustness of an LLM by introducing minor changes to each rule and studying how much the outputs of the original and mutated rules differ.
@@ IV-D3 RQ3
Figure 2 provides examples where, for a given rule and test, both system results should be the same (e.g., Pass, Fail, or NotApplied).
@@ IV-E Statistical Analyses
negligible ,
small ,
medium ,
@@ IV-F Threats to Validity
Internal validity concerns the experiment setup.
We used hosted LLMs, limiting our control over model execution.
While we cannot guarantee that model instances or sessions are not shared across users or HTTP requests, platform documentation suggests each request is treated independently, minimizing leakage risks.
We maintained consistent LLM temperature settings, though results may vary with different values.
To mitigate stochastic effects, we followed best practices, repeating experiments times and using statistical tests [3].
Despite the inherent possibility of bugs in the pipeline, we reviewed scripts and reran experiments to avoid critical errors.
@@
Finally, we perform a laboratory experiment, i.e., we run the tests against a standalone version of GURI in an isolated setting and do not perform a field experiment, where the tests are executed against the real GURI.
V Results and Analyses
V-A RQ1: Effectiveness and Efficiency
@@
Figure 3: Completion rates for all the rules per LLM.
Table II: Statistical tests of the completion rates, success indices per test type, and inference times per LLM
@@
The statistical tests, depicted in Table II, support these observations: GPT-3.5 is statistically better than Llama2 with medium and Mixtral and Mistral with large effect sizes.
Figure 4: Success indices for all the rules per test type and LLM.
@@
This shows that the LLMs struggle more to generate Fail tests, which might be due to the structure of the rules, where the LLM must satisfy the left clause of the implication while it must falsify the right clause (see Section II).
Figure 5: Inference times for all the rules per LLM.
@@
V-B RQ2: Robustness
Figure 6: Robustness indices for all the rules per test type and LLM.
Table III: Statistical tests of the robustness indices for all the rules per test type and LLM
@@ V-B RQ2: Robustness
This is confirmed by the statistical tests (see Table III): GPT-3.5 is better than Mixtral and Llama2 with medium and Mistral with large effect size.
@@
Finally, for , we do not observe statistical differences between Mistral, Mixtral, and GPT-3.5 with %, %, and %, respectively.
@@ V-C RQ3: Differential Testing
A Warning result occurs when the rule execution is successful, but the data exhibits potential issues. For example, a patient’s age might be 120 years, which is theoretically possible but highly unlikely; hence the Warning result. In addition, we identify two other result types:
@@
Table IV shows all the observed matches and mismatches after executing each generated test.
@@
(1) Inconsistency in handling rule versions.We notice that some rules have both versions “active” (e.g., V09/1 and V09/2), while some other rules are “inactive” for the first version. Thus, such rules never Pass.
@@
(3) Inconsistent variable handling.Different variables are handled differently in terms of execution. For instance, “ds” affects the whole cancer message (resulting in empty response), while other variables (e.g., diagnose date) only impact specific rules (always resulting in NotApplied).
Table IV: Matches and mismatches between GURI and Dvare++ results, considering all 58 rules
@@
VI Discussion and Lessons Learned
VI-A Hallucination Effect
As an indicator of potential hallucination, we consider the LLMs accuracy in following the given prompt instructions.
We utilize the metric to determine whether the tests generated by the LLMs are correct.